Multiply and write the result in simplest form. 2/9 x 4/5
Answer by sdmmadam@yahoo.com(530)

Multiply and write the result in simplest form. 2/9 x 4/5
Whenever you multiply two fractions see if there are common elements and if so cancel them. Then multiply the nr with the nr and the dr with the dr.
If there are no common elements straight away multiply the nr with the nr and the dr with the dr.
2/9 x 4/5 = (2X4)/(9X5) = 8/45